OpenStudy (anonymous):

Simplify 5 1/2 + 3/4 - 9/2..

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Do you know about common denominators? @Stephiep75

OpenStudy (skullpatrol):

$$\Huge 5\frac{1}{2}+\frac{3}{4}-\frac{9}{2}$$

OpenStudy (skullpatrol):

If you want to learn @Stephiep75 you have to put in effort.

OpenStudy (anonymous):

first... 5 1/2=10/2+1/2=11/2 then just add the rest with common denominators

OpenStudy (anonymous):

Sorry it wouldn't let me reply... Yes I know common denominators.. Just need some refershing

OpenStudy (anonymous):

Like do I go 5x 2 + 1 = 11/2

OpenStudy (skullpatrol):